OpenStudy (anonymous):

A city has a population of 535,000, and has a growing rate of 4% per year. What is the population after 14 years? round to the nearest whole number.

OpenStudy (anonymous):

@bibby @MTALHAHASSAN2 @shifuyanli

OpenStudy (anonymous):

somebody help me plz

OpenStudy (anonymous):

ill try to help

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

is there any multiple choice answers?

OpenStudy (anonymous):

sadly no :(

OpenStudy (anonymous):

ok, multiply 4% per year by the 14 years

OpenStudy (anonymous):

so times 535,000 by 1.04, then that number by 1.04, and continue on until I've done it 14 times?

OpenStudy (anonymous):

I believe so, just were did you get the 1.04? did you mean 0.04 for 4%?

OpenStudy (radar):

Sounds very similar to an "interest" problem. Where the amount A equals the Principal (initial quantity) times the quatitiy 1 + rate to the power equal to the number of years. For population we could write: P = (initial population) times the quantity 1 + rate of growth to the power equal to the number of years. P = 535,000(1 + .04)^14 Can you solve that? You need to use a calculator with a x^y key.

OpenStudy (radar):

P =535,000 (1.04)^14 P=535,000 X 1.731676448 = 926,447